Please note, the drop down boxes contain A and B which represents the portfolio

Jeanne Lewis is attempting to evaluate two possible portfolios consisting of the same five assets but held in different proportions. She is particularly interested in using beta to compare the risk of the portiolios and. in this regard, has gathered the following data: EEE a. Calculate the betas for potfolios A and B b. Compare the risk of Bach portfolio to the market as well as to each other which portfolio is more risky? a. The beta of portfolio A isRound to three decimal places.) The beta of portiolio B is(Round to three decimal places ) b. Portfolio s slightly less sky than he market average risk while portfolio Portioliois the more risky portfolio. (Select from the drop-down menu) ▼ s more sky than the market Portfolio ▼ s turn will move more than portfolio ▼ for a gl en increase or dec ase ın market risk Select from the drop down menusData Table (Click on the icon located on the top-right corner of the data table in order to copy its contents into a spreadsheet.) Portfolio Weights Asset Asset Beta Portfolio A Portfolio B 1.27 0.73 1.25 1.05 0.92 7% 33% 9% 14% 37% 100% 23% 6% 18% 20% 30% 100% 4 Total Print Done
Answer #1

a). \betap = \sum_{i=1}^{n} [Wi x \betai]

\betaA = [0.07 x 1.27] + [0.33 x 0.73] + [0.09 x 1.25] + [0.14 x 1.05] + [0.37 x 0.92]

= 0.09 + 0.24 + 0.11 + 0.15 + 0.34 = 0.93

\betaB = [0.26 x 1.27] + [0.06 x 0.73] + [0.18 x 1.25] + [0.20 x 1.05] + [0.30 x 0.92]

= 0.33 + 0.04 + 0.23 + 0.21 + 0.28 = 1.09

b). Portfolio B is more risky than the market as well as Portfolio A.

Portfolio A is less risky than the market as well as Portfolio B.